OpenStudy (anonymous):

I need help in 3 Algebra Questions. I will put them in the comment box below <3 :* (Please help will fan & medal if you give me the correct respond) c:

OpenStudy (anonymous):

Use the distributive property to simplify the expression below \[\frac{ 6\times -10 }{ -2}\]

OpenStudy (anonymous):

Use the distributive property to simplify the expression below: \[\frac{ -5\times-10 }{ -5 }\]

OpenStudy (anonymous):

Evaluate: \[3^{2} + (6-2)\times4-\frac{ 6 }{ 3}\]

OpenStudy (anonymous):

6x - 10/-2 6x + 5

OpenStudy (anonymous):

Okay and for questions 2 & 3 ?

OpenStudy (anonymous):

The answer for the evaluating one is: 4x^4+7

OpenStudy (anonymous):

-5x-10/-5 -5x+2

OpenStudy (anonymous):

@Nekojellehfish I don't think so. I have multiple choices here if you would like to see them (:

OpenStudy (anonymous):